OpenStudy (ksaimouli):

an object dropped from rest from the top of a tall building falls y=16t^2 feet first t seconds (a)find teh speed of a object at t=3s and confirm algebraically

OpenStudy (anonymous):

well velocity is just the derivative of the distance equation with respect to time. dy/dt. So the velocty can be determined by 32t evaluated at 3 s= 96 ft/s

OpenStudy (ksaimouli):

hold on i dont know what the derivative is

OpenStudy (ksaimouli):

@Chyme

OpenStudy (anonymous):

are you not in calculus? or possibly pre calc?

OpenStudy (ksaimouli):

i am in cal but now we are learning about limits

OpenStudy (ksaimouli):

we did not start derivatives

OpenStudy (anonymous):

we can do it another way... since we know that velocity is distance over time the distance can be calculated simply by pluging in a time it into y=16t^2 We want to know the velocity at 3 seconds so we need to know the distances at 3 seconds..16*9=144 ft now you just divide this by the time and you have your velocity =144/3 m/s. This is average velocity and the derivative will give you the instaneous velocity which you will get to soon enough.

OpenStudy (ksaimouli):

ok thx

OpenStudy (ksaimouli):

but the answer was 96

OpenStudy (ksaimouli):

where as 144/3 is 48

OpenStudy (anonymous):

It looks like it's asking for instantaneous v.

OpenStudy (anonymous):

So if the initial is zero, and the average is 48, what's the final?

OpenStudy (anonymous):

have you learned anything like this? \[\lim_{\Delta x \rightarrow0 } \frac{ f(x+ \Delta x)-f(x) }{ \Delta x }\]

OpenStudy (anonymous):

my guess is they want him to use the limit definiton of the derivative to figure out the inst velocity

OpenStudy (anonymous):

Maybe, but that's not necessary.

OpenStudy (anonymous):

It would be good exercise and a good intro to derivatives, but this is soluble with basic algebra skills.

OpenStudy (anonymous):

You already found the average velocity by using (total distance)/(total time), but average velocity (for this particular case since the velocity increases linearly) is also (final + initial)/2.

OpenStudy (anonymous):

@ksaimouli, see what i'm saying here? This is an elementary algebra problem. I can show you the calculus necessary to solve it that way, but you don't need it here.